If you flip a coin 3 times what is the probability of getting 3 heads?
- 0.125
- 0.12
- 0.1
- None of these
Explanation
A coin has 2 possible outcomes each time:
- Head (H)
- Tail (T)
If the coin is flipped 3 times, the total possible outcomes are:
2^3 = 8
The possible outcomes are:
HHH, HHT, HTH, HTT, THH, THT, TTH, TTT
Only one outcome gives 3 heads: HHH.
Therefore:
Formula
P(3 heads)=Favourable outcomes/Total outcome=1/8
1/8=0.125 or 12.5%
Related MCQs
- 30
- 25
- 35
- None of these
اس سوال کو وضاحت کے ساتھ پڑھیں
- 41
- 39
- 37
- None of these
اس سوال کو وضاحت کے ساتھ پڑھیں
- 1/9
- 1/6
- 1/12
- None of these
اس سوال کو وضاحت کے ساتھ پڑھیں
- 1/12
- 1/6
- 1/3
- 1/2
اس سوال کو وضاحت کے ساتھ پڑھیں
- 400
- 500
- 600
- 700